OpenStudy (anonymous):

solve -4x+3(x-5)=7 solve for x

OpenStudy (anonymous):

-4x+3x-15 = 7 -x = 7+15 -x = 22 x = -22

OpenStudy (anonymous):

-4x+3x-15=7 -x-15=7 -x=22 x=-22

OpenStudy (anonymous):

awww you beat me to it!